OpenStudy (anonymous):

Solve for x:(using the principle of zero product) 24x^2=8x

OpenStudy (anonymous):

x=1/3

OpenStudy (whpalmer4):

\[24x^2=8x\]\[24x^2-8x = 0\]\[8x(3x-1) = 0\]For that to be true, we must have \[8x=0\rightarrow x = 0\]and\[(3x-1) = 0\rightarrow x=\frac{1}{3}\]